# Rigidity results for initial data sets satisfying the dominant energy condition

### v3

@@ 4. Proof of Theorem I
By Lemma 4.3, we know that
$-\sum_{j=1}^{n-1}\lambda_{j}\,\langle\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j}),s\rangle=\sum_{j=1}^{n-1}\lambda_{j}\,|s|^{2}.$
This implies $-\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j})=s$ whenever $\lambda_{j}>0$. In other words, $\gamma(e_{j})\circ s=s\circ\gamma_{0}(e^{0}_{j})$ whenever $\lambda_{j}>0$. Applying Lemma 4.2 with $X=e_{j}$, we obtain
@@
for $j=1,\ldots,n-1$. By Lemma 4.6, $s$ is invertible at each point in $M$. Consequently,
@@
for $j=1,\ldots,n-1$. Note that $i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$ is self-adjoint, while $\gamma(h(e_{j}))\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ are skew-adjoint. Thus, we conclude that $\gamma(h(e_{j}))=\lambda_{j}\,\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ and $i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$ for $j=1,\ldots,n-1$. From this, the assertion follows. ∎
@@
We now turn to the even-dimensional case of Proposition 4.1. Suppose that $n\geq 2$ is an even integer. Since $\mu\geq|J|$ on $M$ and $H\geq|\mathrm{tr}(q)-q(\nu,\nu)|+\|dN\|_{\mathrm{tr}}$ along $\Sigma$, Corollary 3.12 yields a non-trivial $\nabla^{\mathcal{E},q}$-parallel section $s\in H^{1}(M,\mathcal{E})$. Clearly, $s$ is smooth. Moreover, we can find an element $\hat{t}\in S^{1}$ such that $\gamma(\nu)\circ s=s\circ\gamma_{0}(\hat{N})$ along $\Sigma$. Here, the map $\hat{N}:\Sigma\to S^{n}$ is defined by $\hat{N}(x)=\tilde{N}(x,\hat{t})=\varphi(N(x),\hat{t})$ for all $x\in\Sigma$.
@@
We have
$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}=\|d\hat{N}\|_{\mathrm{tr}}\,|s|^{2}$
at each point on $\Sigma$. Moreover, $\|d\hat{N}\|_{\mathrm{tr}}=\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$.
@@
$=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
This implies
$H\,|s|^{2}+i\,(\mathrm{tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle=-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
Consequently,
$\|dN\|_{\mathrm{tr}}\,|s|^{2}\leq H\,|s|^{2}-|\mathrm{tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}.$
As in Lemma 3.6, we can bound
$-\Big{\langle}\sum_{j=1}^{n-1}\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\Big{\rangle}\leq\|d\hat{N}\|_{\mathrm{tr}}\,|s|^{2}.$
Finally, we note that $\|d\hat{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}$ by definition of $\hat{N}$. Since $|s|^{2}>0$ at each point on $M$, the assertion follows. ∎
Lemma 4.10.
The map $\varphi(\cdot,\hat{t})\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$. In particular, the map $\hat{N}:\Sigma\to S^{n}$ is the composition of the map $N:\Sigma\to S^{n-1}$ with the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
Proof.
Suppose that the assertion is false. By Lemma 3.9, the map $\varphi(\cdot,\hat{t})$ is Lipschitz continuous with a Lipschitz constant strictly less than $1$. This implies $\|d\hat{N}\|_{\mathrm{tr}}<\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$ where $\|dN\|_{\mathrm{tr}}>0$. Using Lemma 4.9, we conclude that $\|dN\|_{\mathrm{tr}}=0$ at each point on $\Sigma$. This implies that $N$ is locally constant. This contradicts the assumption that $N$ has non-zero degree. ∎
@@
$L:=\{\sigma\in S_{0}:\text{\rm$s\sigma=0$ at each point on $M$}\}.$
@@
Since the map $N\colon\Sigma\to S^{n-1}$ has non-zero degree, we know that the image of $N\colon\Sigma\to S^{n-1}$ is all of $S^{n-1}$. Using Lemma 4.10, we conclude that the image of the map $\hat{N}:\Sigma\to S^{n}$ is the equator in $S^{n}$. Thus, we conclude that $\text{\rm span}\{\hat{N}(p):p\in\Sigma\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}$. By Lemma 4.12, $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n}$. Since the volume element acts as a scalar, we deduce that $L$ is invariant under $\gamma_{0}(v)$ for all $v\in\mathbb{R}^{n+1}$. Arguing as in the proof of Lemma 4.6, the assertion follows. ∎
@@ 4.3. Existence of a local isometric immersion into $\mathbb{R}^{n,1}$
Let $\mathcal{T}=TM\oplus\mathbb{R}$ denote the extended tangent bundle. We define a connection $\nabla^{\mathcal{T},q}$ on $\mathcal{T}$ by
@@
for every vector field $X$ and every section $(Y,\varphi)$ of $\mathcal{T}$. We define a bundle metric of signature $(n,1)$ on $\mathcal{T}$ by
for $(Y,\varphi),(Z,\psi)\in\mathcal{T}$.
@@
The connection $\nabla^{\mathcal{T},q}$ is compatible with $\langle\cdot,\cdot\rangle_{\mathcal{T}}$.
@@
Let $X$ be a vector field, and let $(Y,\varphi)$, $(Z,\psi)$ be two sections of $\mathcal{T}$. Then
@@
$=g(D_{X}Y,Z)+g(Y,D_{X}Z)-X(\varphi)\,\psi-\varphi\,X(\psi)$
$=X(g(Y,Z)-\varphi\psi)$
$=X(\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}).$
@@
Suppose that $u,v$ are $\nabla^{\mathcal{S},q}$-parallel sections of $\mathcal{S}$. We define a vector field $Y$ on $M$ by the requirement $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ for all $Z\in TM$. Moreover, we define a scalar function $\varphi$ on $M$ by $\varphi=-\text{\rm Re}(\langle u,v\rangle)$. Then $(Y,\varphi)$ is a $\nabla^{\mathcal{T},q}$-parallel section of $\mathcal{T}$.
@@
$=-\text{\rm Re}(\langle u,v\rangle)\,q(X,Z)$
$=\varphi\,q(X,Z)$
@@
Let $p$ be an arbitrary point in $M$. Consider the set of all elements $(Y,\varphi)\in\mathcal{T}|_{p}$ which can be written in the form $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ and $\varphi=-\text{\rm Re}(\langle u,v\rangle)$, where $u,v\in\mathcal{S}|_{p}$. This set spans $\mathcal{T}|_{p}$.
@@
Suppose that the assertion is false. Then there exists a non-zero element $(Z,\psi)\in\mathcal{T}|_{p}$ such that
@@
for all $u,v\in\mathcal{S}|_{p}$. From this, we deduce that $i\,\gamma(Z)+\psi\,\text{\rm id}=0\in\text{\rm End}(\mathcal{S}|_{p})$. This implies $Z=0$ and $\psi=0$. This is a contradiction. ∎
@@
By Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$. Consequently, the spinor bundle $\mathcal{S}$ admits a trivialization by $\nabla^{\mathcal{S},q}$-parallel sections. Using Lemma 4.17, we obtain a collection of $\nabla^{\mathcal{T},q}$-parallel sections of $\mathcal{T}$. At each point $p\in M$, these $\nabla^{\mathcal{T},q}$-parallel sections span the fiber $\mathcal{T}|_{p}$ (see Lemma 4.18). This completes the proof of Proposition 4.19. ∎
@@
where $m_{\alpha\beta}$ is a constant $(n+1)\times(n+1)$-matrix. The matrix $m_{\alpha\beta}$ is invertible with signature $(n,1)$. By taking suitable linear combinations, we may arrange that
@@
(4.4) $00=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,g(E^{(\alpha)},Y)\,\zeta^{(\beta)}=$ 0
@@
(4.5) $-1=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,\zeta^{(\alpha)}\,\zeta^{(\beta)}.$
Since $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it follows that the covariant derivative satisfies $DE^{(\alpha)}=\zeta^{(\alpha)}q$ and is hence symmetric. In particular, the one-form $g(E^{(\alpha)},\cdot)$ is closed. We may locally write $g(E^{(\alpha)},\cdot)=df^{(\alpha)}$ for some scalar function $f^{(\alpha)}$. It follows from (4.3) that $g=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,df^{(\alpha)}\otimes df^{(\beta)}$. In other words, the map $F=(f^{(0)},f^{(1)},\ldots,f^{(n)})$ is a local isometric immersion of $(M,g)$ into the Minkowski spacetime $\mathbb{R}^{n,1}$. For each vector $X\in TM$, we have
@@
It follows from (4.4) and (4.5) that the unit normal vector to the immersion $F$ is given by $(\zeta^{(0)},\zeta^{(1)},\ldots,\zeta^{(n)})\in\mathbb{R}^{n,1}$. Using the fact that $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it is easy to see that the second fundamental form associated with $F$ is equal to $q$.
Note that the isometric immersion $F=(f^{(0)},f^{(1)},\ldots,f^{(n)})$ is indeed globally well-defined if $M$ is simply connected. If $M$ is not simply connected, then $F$ may not be globally well-defined, but its differential is globally well-defined.

### v4

@@ 4. Proof of Theorem I
Applying Lemma 4.2 with $X=e_{j}$, we obtain
(4.3) $\gamma(h(e_{j}))\circ s+i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\circ s=\lambda_{j}\,s\circ\gamma_{0}(e^{0}_{j})$
for $j=1,\ldots,n-1$. This implies
$H\,$ $|s|^{2}+i\,(\text{\rm tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle$
$=-\sum_{j}\langle\gamma(e_{j})\circ\gamma(h(e_{j}))\circ s,s\rangle-i\sum_{j}\,\langle\gamma(e_{j})\circ\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\circ s,s\rangle$
$=-\sum_{j}\lambda_{j}\,\langle\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j}),s\rangle.$
In the next step, we take the absolute value on both sides. Using the triangle inequality, we obtain
(4.4) $H\,|s|^{2}-|\text{\rm tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq\|dN\|_{\mathrm{tr}}\,|s|^{2}.$
On the other hand, our assumption implies that $\|dN\|_{\mathrm{tr}}\leq H-|\text{\rm tr}(q)-q(\nu,\nu)|$. Consequently, equality holds in (4.4). From this, we deduce that $\gamma(e_{j})\circ s\circ\gamma_{0}(e^{0}_{j})=s$ whenever $\lambda_{j}>0$. Substituting this into (4.3) gives
@@
for $j=1,\ldots,n-1$. Since $s\in\text{\rm Hom}(S_{0},\mathcal{S}|_{p})$ is invertible by Lemma 4.5, it follows that
@@
for $j=1,\ldots,n-1$. Note that
$i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\in\text{\rm End}(\mathcal{S}|_{p})$
is self-adjoint, while $\gamma(h(e_{j}))\in\text{\rm End}(\mathcal{S}|_{p})$ and $\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ are skew-adjoint. Thus, we conclude that $\gamma(h(e_{j}))=\lambda_{j}\,\gamma(e_{j})\in\text{\rm End}(\mathcal{S}|_{p})$ and $i\,\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)=0\in\text{\rm End}(\mathcal{S}|_{p})$ for $j=1,\ldots,n-1$. From this, the assertion follows. ∎
@@
[…]
for all $x\in\Sigma$.
@@
We have $\|d\hat{N}\|_{\mathrm{tr}}=\|dN\|_{\mathrm{tr}}$ at each point on $\Sigma$.
@@
$H\,$ $|s|^{2}+i\,(\text{\rm tr}(q)-q(\nu,\nu))\,\langle\gamma(\nu)\circ s,s\rangle$
$=-\sum_{j}\langle\gamma(e_{j})\circ\gamma(h(e_{j}))\circ s,s\rangle-i\sum_{j}\,\langle\gamma(e_{j})\circ\gamma(q(e_{j})^{\text{\rm tan}})\circ\gamma(\nu)\circ s,s\rangle$
$=-\sum_{j}\langle\gamma(e_{j})\circ s\circ\gamma_{0}(d\hat{N}(e_{j})),s\rangle.$
In the next step, we take the absolute value on both sides. This gives
$H\,|s|^{2}-|\text{\rm tr}(q)-q(\nu,\nu)|\,|s|^{2}\leq\|d\hat{N}\|_{\mathrm{tr}}\,|s|^{2}.$
Since $|s|^{2}>0$ at each point in $M$, it follows that $H-|\text{\rm tr}(q)-q(\nu,\nu)|\leq\|d\hat{N}\|_{\mathrm{tr}}$. On the other hand, $\|d\hat{N}\|_{\mathrm{tr}}\leq\|dN\|_{\mathrm{tr}}$ by definition of $\hat{N}$. Finally, our assumption implies that
$\|dN\|_{\mathrm{tr}}\leq H-|\text{\rm tr}(q)-q(\nu,\nu)|.$
Putting these facts together, the assertion follows. ∎
Lemma 4.9.
The map $\varphi(\cdot,\hat{t})\colon S^{n-1}\to S^{n}$ is the standard inclusion $S^{n-1}\hookrightarrow S^{n}$. In particular, the map $\hat{N}\colon\Sigma\to S^{n}$ is the composition of the map $N\colon\Sigma\to S^{n-1}$ with the standard inclusion $S^{n-1}\hookrightarrow S^{n}$.
Proof.
@@
$L:=\{\sigma\in S_{0}\colon\text{\rm$s\sigma=0$ at each point in $M$}\}.$
@@
Since the map $N\colon\Sigma\to S^{n-1}$ has non-zero degree, we know that the image of $N\colon\Sigma\to S^{n-1}$ is all of $S^{n-1}$. Using Lemma 4.9, we conclude that the image of the map $\hat{N}\colon\Sigma\to S^{n}$ is the equator in $S^{n}$. In particular, $\text{\rm span}\{\hat{N}(p)\colon p\in\Sigma\}=\mathbb{R}^{n}\subset\mathbb{R}^{n+1}$. By Lemma 4.11, $L$ is invariant under $\gamma_{0}(\xi)$ for all $\xi\in\mathbb{R}^{n}$. Since the volume element acts as a scalar multiple of the identity, we deduce that $L$ is invariant under $\gamma_{0}(\xi)$ for all $\xi\in\mathbb{R}^{n+1}$. Arguing as in the proof of Lemma 4.5, the assertion follows. ∎
@@ 4.3. Existence of a local isometric immersion into $\mathbb{R}^{n,1}$
We define a vector bundle $\mathcal{T}$ over $M$ by $\mathcal{T}=TM\oplus\mathbb{R}$. We define a bundle metric of signature $(n,1)$ on $\mathcal{T}$ by
for $(Y,\varphi),(Z,\psi)\in\mathcal{T}$. We define a connection $\nabla^{\mathcal{T},q}$ on $\mathcal{T}$ by
@@
for every vector field $X$ and every section $(Y,\varphi)$ of $\mathcal{T}$.
@@
The connection $\nabla^{\mathcal{T},q}$ is compatible with the bundle metric $\langle\cdot,\cdot\rangle_{\mathcal{T}}$.
@@
Let $X$ be a vector field on $M$, and let $(Y,\varphi)$ and $(Z,\psi)$ be two sections of $\mathcal{T}$. Then
$\langle\nabla_{X}^{\mathcal{T},q}(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}=g(D_{X}Y,Z)-\varphi\,q(X,Z)-X(\varphi)\,\psi+q(X,Y)\,\psi$
and
$\langle(Y,\varphi),\nabla_{X}^{\mathcal{T},q}(Z,\psi)\rangle_{\mathcal{T}}=g(Y,D_{X}Z)-q(X,Y)\,\psi-\varphi\,X(\psi)+\varphi\,q(X,Z).$
This gives
@@
$=X(g(Y,Z)-\varphi\psi)=X(\langle(Y,\varphi),(Z,\psi)\rangle_{\mathcal{T}}).$
@@
Suppose that $u,v$ are $\nabla^{\mathcal{S},q}$-parallel sections of $\mathcal{S}$. We define a vector field $Y$ on $M$ so that $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ for all $Z\in TM$. Moreover, we define a scalar function $\varphi$ on $M$ by $\varphi=-\text{\rm Re}(\langle u,v\rangle)$. Then $(Y,\varphi)$ is a $\nabla^{\mathcal{T},q}$-parallel section of $\mathcal{T}$.
@@
$=-q(X,Z)\,\text{\rm Re}(\langle u,v\rangle)$
$=q(X,Z)\,\varphi$
@@
Suppose that $p$ is an arbitrary point in $M$. Consider the set of all pairs $(Y,\varphi)\in\mathcal{T}|_{p}$ with the property that there exist $u,v\in\mathcal{S}|_{p}$ such that $g(Y,Z)=\text{\rm Re}(i\,\langle\gamma(Z)\,u,v\rangle)$ for all $Z\in TM$ and $\varphi=-\text{\rm Re}(\langle u,v\rangle)$. This set spans $\mathcal{T}|_{p}$.
@@
Suppose that the assertion is false. Then there exists a non-zero pair $(Z,\psi)\in\mathcal{T}|_{p}$ such that
@@
for all $u,v\in\mathcal{S}|_{p}$. This implies that $i\,\gamma(Z)+\psi\,\textup{id}=0\in\text{\rm End}(\mathcal{S}|_{p})$. From this, we deduce that $Z=0$ and $\psi=0$. This is a contradiction. ∎
@@
According to Proposition 4.1, there exists a $\nabla^{\mathcal{E},q}$-parallel section $s$ of the bundle $\text{\rm Hom}(S_{0},\mathcal{S})$ with the property that $s$ is invertible at each point in $M$. Consequently, the spinor bundle $\mathcal{S}$ admits a trivialization by $\nabla^{\mathcal{S},q}$-parallel sections. Using Lemma 4.16, we obtain a collection of $\nabla^{\mathcal{T},q}$-parallel sections of $\mathcal{T}$. At each point $p\in M$, these $\nabla^{\mathcal{T},q}$-parallel sections span the fiber $\mathcal{T}|_{p}$ (see Lemma 4.17). This completes the proof of Proposition 4.18. ∎
@@
where $m_{\alpha\beta}$ is a constant $(n+1)\times(n+1)$-matrix. The matrix $m_{\alpha\beta}$ is invertible and has signature $(n,1)$. By taking suitable linear combinations, we may arrange that
@@
(4.6) $0=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,g(E^{(\alpha)},Y)\,\zeta^{(\beta)}=0$
@@
(4.7) $1=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,\zeta^{(\alpha)}\,\zeta^{(\beta)}.$
Since $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it follows that the covariant derivative of $E^{(\alpha)}$ is given by $\zeta^{(\alpha)}q$ and is hence symmetric. Thus, the one-form $g(E^{(\alpha)},\cdot)$ is closed. We may locally write $g(E^{(\alpha)},\cdot)=df^{(\alpha)}$ for some scalar function $f^{(\alpha)}$. It follows from (4.5) that $g=\sum_{\alpha,\beta=0}^{n}m_{\alpha\beta}\,df^{(\alpha)}\otimes df^{(\beta)}$. In other words, the map $F=(f^{(0)},f^{(1)},\ldots,f^{(n)})$ is a local isometric immersion of $(M,g)$ into the Minkowski spacetime $\mathbb{R}^{n,1}$. For each vector $X\in TM$, we have
@@
It follows from (4.6) and (4.7) that the unit normal vector to the immersion $F$ in $\mathbb{R}^{n,1}$ is given by $(\zeta^{(0)},\zeta^{(1)},\ldots,\zeta^{(n)})$. Since $(E^{(\alpha)},\zeta^{(\alpha)})$ is $\nabla^{\mathcal{T},q}$-parallel, it is easy to see that the second fundamental form associated with the immersion $F$ is equal to $q$.
If $M$ is simply connected, the isometric immersion $F=(f^{(0)},f^{(1)},\ldots,f^{(n)})$ is globally well-defined. If $M$ is not simply connected, then $F$ may not be globally well-defined, but its differential is globally well-defined.